OpenStudy (snowcrystal):

Help!!! I need someone to explain step by step how to do this problem. I will fan and medal

OpenStudy (snowcrystal):

Factor completely 48y^2 + 82y -70

OpenStudy (lukebluefive):

First, write out the problem: \[48y^2 + 82y - 70\] Then, remove the common factor: \[2*(24y^2 + 41y - 35)\]

OpenStudy (lukebluefive):

Now, focus on the part in parenthesis: \[24y^2 + 41y - 35\] In order to factor this, it needs to be written in the following form: \[(ay + c)(by - d)\] Where a, b, c, and d are constants.

OpenStudy (lukebluefive):

Next, we list the products of 24 to find the possibilities of a and b: 24 = 2 * 12 = 3 * 8 = 4 * 6 So a and b could equal 2 and 12, 3 and 8, or 4 and 6 (not necessarily respectively).

OpenStudy (lukebluefive):

Do the same to 35 to find the possible values for c and d: 35 = 5 * 7 So either c = 5 and d = 7, or c = 7 and d = 5.

OpenStudy (lukebluefive):

Sorry, it's -35, so this is the proper product listing: 35 = -5 * 7 = 5 * -7 So c and d could equal -5 and 7, or 5 and -7 (again, not necessarily respectively).

OpenStudy (lukebluefive):

Since the middle term (41y) is positive, we'll guess that c = -5 and d = 7, since if the larger number were negative, the middle term would most likely be negative too. This leaves us with: \[(ay - 5)(by + 7)\]

OpenStudy (lukebluefive):

At this point, we try combinations of a and b until it fits the given value: a = 2, b = 12 \[(2y - 5)(12y + 7) = 24y^2 - 60y + 14y - 35 = 24y^2 - 46y - 35\] a = 12, b = 2 \[(12y - 5)(2y + 7) = 24y^2 + 84y - 10y - 35 = 24y^2 + 74y - 35\] a = 3, b = 8 \[(3y - 5)(8y + 7) = 24y^2 - 40y + 21y - 35 = 24y^2 - 19y - 35\] a = 8, b = 3 \[(8y - 5)(3y + 7) = 24y^2 - 15y + 56y - 35 = 24y^2 + 41y - 35\]

OpenStudy (snowcrystal):

it would be a=8 , b=3 right?

OpenStudy (lukebluefive):

Yes, so: \[24y^2 + 41y - 35 = (8y - 5)(3y+7)\] Those are its factors.